# From Kernels to Attention: A Transformer Framework for Density and Score Estimation

### v1

@@ 4.1 Score estimation
To compare our model to KDE, we first sample a 2d Gaussian, visualize the estimated scores using a quiver plot and compute the MSE score loss. Figure 3 demonstrates that even in dimension 2 and KDE-favorable example of a Gaussian, our model performs better. The KDE score approximation is computed using
$k(x_{i},x_{j})$ $=\exp\!\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2h^{2}}\right),$ (8)
$\widehat{s}(x_{i})$ $=\frac{1}{h^{2}}\left(\frac{\sum_{j}k(x_{i},x_{j})x_{j}}{\sum_{j}k(x_{i},x_{j})}-x_{i}\right).$ (9)
Figure 3: Score estimation comparison between Silverman KDE (Silverman, 2018) and our transformer model. The transformer is more accurate, especially in the sparse regions. We plot the negated score for easier viewing.
For a quantitative comparison of score estimation against the KDE, we compute the Mean Squared Error (MSE) for dimensions $1$ and $10$ across several magnitudes of samples sizes $n$. See the left panel of Figure 4. The transformer model achieves much better accuracy than KDE even in dimension 1 and especially in dimension 10.
To judge the out-of-distribution generalization of our model, we compare the MSE loss on GMMs with 1-19 modes, while the training data consists of GMMs with 1-10 modes. The right panel of Figure 4 shows that the MSE loss is monotone and stable in the number of modes. This indicates good generalization. Additionally, we evaluate on another type of distribution, the Laplace distribution, which is more peaked near the origin. Despite being only trained on GMMs, the model is able to estimate the score successfully, as shown in table 1.
Table 1: MSE loss on the 2D Laplace distribution.
$n$ Transformer KDE
512 0.3598 0.3810
1024 0.2992 0.3305
2048 0.2756 0.2990
4096 0.2597 0.2650
@@ 4.3 Density Estimation via SD-KDE
which counteracts KDE’s smoothing bias. In practice the true score is unavailable. We therefore consider three ways to obtain it: (i) Empirical SD-KDE (Emp-SD-KDE), which uses the Silverman KDE to approximate the score; (ii) SD-KDE using our Transformer’s predicted score; and (iii) SD-KDE using the autograd score of a Transformer-learned log-density. For completeness, we compare these against (iv) plain KDE with Silverman bandwidth (Silverman, 2018) and (v) the Transformer’s direct density/log-density output.
@@
    Emp-SD-KDE: SD-KDE with the score derived from Silverman KDE.
@@
    SD-KDE (Autograd Score): SD-KDE with the score from autograd of the Transformer’s learned log-density (Paszke et al., 2019).
@@
Figure 6 shows that incorporating a learned score substantially improves density recovery over plain KDE; Emp-SD-KDE already reduces bias, while SD-KDE with our predicted score yields the most accurate estimate. In 1D bimodal GMMs (Figure 7), SD-KDE with the learned score best matches the truth by MSE. Across sample sizes (Figure 8), Emp-SD-KDE converges faster than KDE, but our Transformer-based approaches exhibit superior accuracy and scaling. The direct density approach performs best at small $n$ but, as expected, scaling with $n$ favors score-based methods; learning the score is an easier objective than learning the density itself. Notably, our Transformer was trained only at $n=2048$ yet generalizes well to other sample sizes.
Figure 8: MSE vs. sample size in 1D (left) and 2D (right). Emp-SD-KDE improves over KDE; SD-KDE with our learned score and the Transformer model show the best scaling. The Transformer was trained only at $n=2048$.
@@ 4.4 Estimation of entropy and Fisher information
Differential entropy $H(f)=-\mathbb{E}_{f}\log f$ and its gradient, Fisher information $I(f)=\mathbb{E}_{f}\|\nabla\log f\|_{2}^{2}$, are fundamental statistical quantities that characterize the spread of the distribution. If we use the discretizations
@@
to estimate $H(f)$ and $I(f)$ using only an iid sample of $f$. In dimension 1 we sample a GMM with three modes, numerically compute its entropy and Fisher information using expensive grid integration and the oracle knowledge of its density and score, and compute the estimation error. In dimension 10 we cannot afford to do grid estimation, so we sample a random gaussian with randomized mean and covariance, and use the known analytical formulas. We use the debiased KDE method of (Epstein et al., 2025) here as well. Figure 9 demonstrates that already in dimension 1 both the score model $S$ and the density model $T$ outperform Silverman KDE. In dimension 10 the learned density performs much better for entropy estimation and and the learned score performs much better for Fisher information estimation.
Figure 9: Comparison between the transformer model (learned) and Silverman KDE and score-debiased KDE for estimation of differential entropy $H(f)$ and Fisher Information $I(f)$. Transformer’s MSE error is lower than the error of the KDE approximation, even in dimension 1.
@@ 4.5 Plasma simulation
$v(x)$ $=-\int_{\mathbb{R}^{3}}A(x-y)(\nabla\log f(x)-\nabla\log f(y))f(y)dy$ (12)
$A(z)$ $=\|z\|_{2}^{\gamma}\left(\|z\|_{2}^{2}I_{3}-zz^{T}\right)$ (13)
and solved numerically with a particle method by approximating the score via a particular KDE variant, called the blob method. On the other hand, in machine learning Boffi & Vanden-Eijnden (2023) proposed a NN-based method of solving any Fokker-Planck equation called Score Based Transport Modeling, while Ilin et al. (2025a, b) adapted it to the Landau equation and the problem of sampling. The core idea is to continually train the NN $s_{\theta}$ on the score matching loss $\mathbb{E}_{f}\left[\|s_{\theta}(x)\|^{2}+2\nabla_{x}\cdot s_{\theta}(x)\right]$, which is equivalent to training on the MSE loss $\mathbb{E}_{f}\left[\|s_{\theta}-\nabla\log f\|^{2}\right]$. While Boffi & Vanden-Eijnden (2023); Ilin et al. (2025a); Ilin (2025) proved theoretically and demonstrated empirically that this online training effectively bounds the error, using SBTM can be slow due to the on-the-fly NN training. We reproduce the experiments 5.3 and 5.4 from (Ilin et al., 2025a) using the learned score mapping $S(X)\mapsto\nabla\log f(X)$. Unlike SBTM, we only use the pre-trained model and do not perform any training throughout the simulation. Figure 10 demonstrates that the Transformer model is able to match the analytic covariance well, while the KDE-based solver struggles.
Figure 10: Comparison between the trained Transformer model and Silverman KDE on the task of numerically solving the homogeneous Landua equation. We plot the first entry of the covariance matrix $\Sigma_{1,1}(t)$ of the numerical solutions and ground truth, when known. The left two panels use Maxwell collisions, $\gamma=0$ in (13), while the last panel shows Coulomb collisions, $\gamma=-3$. The Transformer performs better than KDE, and is comparable in quality to SBTM in (Ilin et al., 2025a).
4.6 Runtime
While KDE is known to fail in moderate and high dimensions, it is thought to be very fast. We compare the wall-clock runtime of KDE score estimation as in equation (9), and the Transformer model we propose. The Transformer model contains $800,000$ learned parameters. We use both KDE and the Transformer model on a single l40s GPU with 48Gb of memory. Figure 11 shows that while KDE is faster for small sample size $n\leq 2048$, it becomes slower after this point, especially in higher dimension. The Transformer exhibits sub-quadratic runtime scaling, whereas the KDE has quadratic scaling in dimension 2 and $O(n^{2.5})$ scaling in dimension 10. Additionally, the KDE runs out of GPU memory after $n=32,768$.
Figure 11: The runtime comparison between KDE and the transformer model in 2 and 10 dimensions. KDE encounters an OOM error at $n=2^{15}$. The transformer model exhibits better runtime and memory scaling both in sample size $n$ and dimension $d$.
@@ 5 Conclusion
We present a unified attention-based framework for estimating both the density and score of a distribution from samples. By combining cross-attention with lightweight symmetry handling, the model learns to infer values at arbitrary query points while respecting basic invariances such as permutation and affine transformations. A single network jointly predicts the density and its score, adapting seamlessly across sample sizes and dimensions. Experiments show that this approach achieves consistent accuracy and scaling advantages over classical estimators, bridging the gap between traditional nonparametric density estimation and modern deep learning methods.
@@ Appendix A Proofs
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, and $\mu\in\mathbb{R}^{d}$. Then
$T\big(PXA+\mu\big)$ $=P\,|\det A|^{-1}\,T(X),$ (14)
$S\big(PXA+\mu\big)$ $=P\,(A^{T})^{-1}\,S(X).$ (15)
@@
Consider the density obtained from $f$ by the affine change of variables $y=Ax+\mu$:
$f^{(\mu,A)}(y):=|\det A|^{-1}f\big(A^{-1}(y-\mu)\big).$
@@
$\nabla_{y}\log f^{(\mu,A)}(y)=\big(A^{-1}\big)^{T}\nabla_{x}\log f(x)\big|_{x=A^{-1}(y-\mu)}=(A^{T})^{-1}\nabla\log f\big(A^{-1}(y-\mu)\big).$
Evaluating at $y_{i}=Ax_{\sigma(i)}+\mu$ gives
$\nabla_{y}\log f^{(\mu,A)}(y_{i})=(A^{T})^{-1}\nabla\log f(x_{\sigma(i)}).$
@@
$S(Y)=\begin{pmatrix}\nabla_{y}\log f^{(\mu,A)}(y_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla_{y}\log f^{(\mu,A)}(y_{n})\end{pmatrix}=\begin{pmatrix}(A^{T})^{-1}\nabla\log f(x_{\sigma(1)})\\[-2.0pt] \vdots\\[-2.0pt] (A^{T})^{-1}\nabla\log f(x_{\sigma(n)})\end{pmatrix}=P\,(A^{T})^{-1}\,S(X),$
@@
Proposition A.2 (Attention computes KDE).
@@ Appendix B Attention visualization
To interpret the outputs of the Transformer model we visualize the attention of the eight individual heads in Layer 0 in two different ways. First, in Figure 12 we show the full attention matrices as heatmaps. We choose the particle ordering so that nearby particles are close in the ordering. Second, in Figure 13 we choose a random query point, marked with a red x and color other points according to the strength of the query point’s attention. We clearly observe an emergent behavior – heads specialize in different tasks. Head 1 specialized to look at far-away points, whereas Heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, Heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with multi-scale learned kernels.
@@ Appendix C Comparison with the Score Matching Loss
which corresponds to the Fisher divergence $\mathbb{E}_{p}[\|s_{\theta}(x)-\nabla_{x}\log p(x)\|^{2}]$ up to a constant.

### v2

@@ 4.1 Score estimation
Figure 4: MSE of score estimation on the Laplace distribution. Test-time training (TTT) improves the out-of-distribution generalization.
To compare our model to KDE, we first sample a 2d Gaussian, visualize the estimated scores using a quiver plot and compute the MSE score loss. Figure 3 demonstrates that even in dimension 2 and a KDE-favorable example of a unimodal Gaussian, our model performs better. The KDE score approximation is computed using
$k(x_{i},x_{j})$ $=\exp\!\left(-\frac{1}{2}\sum_{\ell=1}^{d}\frac{(x_{i}^{\ell}-x_{j}^{\ell})^{2}}{h_{\ell}^{2}}\right),$ (9)
$\widehat{s}(x_{i})$ $=\frac{1}{h^{2}}\odot\left(\frac{\sum_{j}k(x_{i},x_{j})x_{j}}{\sum_{j}k(x_{i},x_{j})}-x_{i}\right),$ (10)
where $h_{\ell}=\sigma_{\ell}n^{-1/(d+4)}$ is Silverman’s rule (Silverman, 1986) applied per coordinate with $\sigma_{\ell}$ the sample standard deviation along dimension $\ell$, and $\odot$ denotes element-wise multiplication.
For a quantitative comparison of score estimation against the KDE, we compute the Mean Squared Error (MSE) for dimensions $1$ and $10$ across several magnitudes of sample sizes $n$. See the left panel of Figure 5. The transformer model achieves much better accuracy than KDE even in dimension 1 and especially in dimension 10.
To judge the out-of-distribution generalization of our model, we compare the MSE loss on GMMs with 1-19 modes, while the training data consists of GMMs with 1-10 modes. The right panel of Figure 5 shows that the MSE loss is monotone and stable in the number of modes. This indicates good generalization. Additionally, we evaluate on two non-Gaussian distributions: the Laplace distribution, which has a sharper peak and heavier tails than a Gaussian, and the Student-$t$ distribution ($\nu=3$), which has polynomial rather than exponential tail decay. Despite being trained only on GMMs, the model estimates the score successfully on both, as shown in Tables 2 and 3. Additionally, we can further improve the out-of-distribution performance using test-time training (TTT) (Sun et al., 2020; Gandelsman et al., 2022), using the consistency loss
$\mathcal{L}_{\text{con}}=\frac{1}{n}\sum_{i=1}^{n}\left\|S(X)_{i}-\nabla_{x_{i}}\log T(X)_{i}\right\|_{2}^{2},$ (11)
where $T(X)_{i}$ is the predicted density and $\nabla\log T(X)$ is computed via automatic differentiation. Figure 4 and Table 3 show that as few as 4 TTT steps are sufficient to improve performance on both distributions.
Table 2: MSE of score estimation on the 2D Laplace distribution.
$n$ KDE Transformer
512 0.3810 0.3598
1024 0.3305 0.2992
2048 0.2990 0.2756
4096 0.2650 0.2597
Table 3: MSE of score estimation on the 2D Student-$t$ distribution ($\nu=3$).
$n$ KDE No TTT TTT 4 TTT 6 TTT 8
128 0.1515 0.1980 0.1676 0.1569 0.1450
256 0.1206 0.1119 0.0956 0.0908 0.0897
512 0.0916 0.0574 0.0488 0.0485 0.0514
1024 0.0812 0.1023 0.0806 0.0768 0.0765
@@ 4.3 Density Estimation via SD-KDE
Figure 9: MSE of density estimation in 1D (left) and 2D (right). SD-KDE with our learned score and the Transformer model show the best scaling. The Transformer was trained only at $n=2048$.
@@
which counteracts KDE’s smoothing bias. In practice the true score is unavailable. We therefore consider five methods: (i) plain KDE with Silverman bandwidth (Silverman, 1986); (ii) Empirical SD-KDE (Emp-SD-KDE), which uses the Silverman KDE to approximate the score; (iii) SD-KDE using our Transformer’s predicted score; (iv) SD-KDE using the autograd score of a Transformer-learned log-density; and (v) the Transformer’s direct density/log-density output.
@@
    Emp-SD-KDE: SD-KDE with the score derived from Silverman KDE, following (Epstein et al., 2025).
@@
    SD-KDE (Autograd Score): SD-KDE with the score from autograd of the Transformer’s learned log-density.
@@
Figure 7 shows that incorporating a learned score substantially improves density recovery over plain KDE; Emp-SD-KDE reduces bias, while SD-KDE with our predicted score yields the most accurate estimate. In 1D bimodal GMMs (Figure 8), SD-KDE with the learned score best matches the truth by MSE. Our Transformer-based approaches exhibit superior accuracy and scaling. The direct density approach performs best at small $n$ but, as expected, scaling with $n$ favors score-based methods; learning the score is an easier objective than learning the density itself due to the difficulty of learning the normalization constant $Z=\int_{\mathbb{R}^{d}}f$.
@@ 4.4 Estimation of entropy and Fisher information
Differential entropy $H(f)=-\mathbb{E}_{f}\log f$ and its gradient, Fisher information $I(f)=\mathbb{E}_{f}\|\nabla\log f\|_{2}^{2}$, are fundamental statistical quantities that characterize the spread of the distribution. We use the discretizations
@@
to estimate $H(f)$ and $I(f)$ using only an iid sample of $f$. In dimension 1 we sample a GMM with three modes, numerically compute its entropy and Fisher information using expensive grid integration and the oracle knowledge of its density and score, and compute the estimation error. In dimension 10 we cannot afford to do grid estimation, so we sample a random Gaussian with randomized mean and covariance, and use the known analytical formulas. We use the debiased KDE method of (Epstein et al., 2025) here as well. Figure 10 demonstrates that already in dimension 1 both the score model $S$ and the density model $T$ outperform Silverman KDE. In dimension 10 the learned density performs much better for entropy estimation and the learned score performs much better for Fisher information estimation.
Figure 10: Comparison between the transformer model (learned) and Silverman KDE and score-debiased KDE for estimation of differential entropy $H(f)$ and Fisher Information $I(f)$. Transformer’s MSE is lower than that of the KDE approximation, even in dimension 1.
@@ 4.5 Plasma simulation
$v(x)=-\int_{\mathbb{R}^{3}}A(x-y)(\nabla\log f(x)-\nabla\log f(y))f(y)dy$
and solved numerically with a particle method by approximating the score via a particular KDE variant, called the blob method. On the other hand, in machine learning Boffi and Vanden-Eijnden (2023) proposed a NN-based method of solving any Fokker-Planck equation called Score Based Transport Modeling, while Ilin et al. (2025a, b) adapted it to the Landau equation and the problem of sampling. The core idea is to continually train the NN $s_{\theta}$ on the score matching loss $\mathbb{E}_{f}\left[\|s_{\theta}(x)\|^{2}+2\nabla_{x}\cdot s_{\theta}(x)\right]$, which is equivalent to training on the MSE loss $\mathbb{E}_{f}\left[\|s_{\theta}-\nabla\log f\|^{2}\right]$. While Boffi and Vanden-Eijnden (2023); Ilin et al. (2025a); Ilin (2025) proved theoretically and demonstrated empirically that this online training effectively bounds the error, using SBTM can be slow due to the on-the-fly NN training. We reproduce the experiments 5.3 and 5.4 from (Ilin et al., 2025a) using the learned score. Unlike SBTM, we only use the pre-trained model and do not perform any training throughout the simulation. Figure 11 demonstrates that the Transformer model is able to match the analytic covariance well, while the KDE-based solver struggles.
Figure 11: Comparison between the trained Transformer model and Silverman KDE on the task of numerically solving the homogeneous Landau equation. We plot the first entry of the covariance matrix $\Sigma_{1,1}(t)$ of the numerical solutions and ground truth, when known. The left two panels use Maxwell collisions, while the last panel shows Coulomb collisions. The Transformer outperforms KDE, and is comparable in quality to SBTM in (Ilin et al., 2025a).
@@ 5 Conclusion
We introduced DiScoFormer, a permutation- and affine-equivariant Transformer that jointly estimates density and score from i.i.d. samples in a single forward pass. We proved that self-attention recovers normalized KDE and confirmed empirically that individual heads learn multi-scale kernel-like behaviors. DiScoFormer outperforms KDE across dimensions and sample sizes and serves as a plug-in oracle for score-debiased KDE, entropy and Fisher information estimation, and Fokker-Planck-type PDEs.
@@ Appendix A Proofs
Let $P\in\mathbb{R}^{n\times n}$ be a permutation matrix, $A\in\mathbb{R}^{d\times d}$ be invertible, $\mu\in\mathbb{R}^{d}$, and $\mathbf{1}\in\mathbb{R}^{n}$ be the vector of ones. Then
$T\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,|\det A|^{-1}\,T(X),$ (14)
$S\big(PXA+\mathbf{1}\mu^{\top}\big)$ $=P\,A^{-1}\,S(X).$ (15)
@@
Consider the density obtained from $f$ by the affine change of variables $y=A^{\top}x+\mu$:
$f^{(\mu,A)}(y):=|\det A|^{-1}f\big((A^{\top})^{-1}(y-\mu)\big).$
@@
$\nabla_{y}\log f^{(\mu,A)}(y)=\big((A^{\top})^{-1}\big)^{\top}\nabla_{x}\log f(x)\big|_{x=(A^{\top})^{-1}(y-\mu)}=A^{-1}\nabla\log f\big((A^{\top})^{-1}(y-\mu)\big).$
Evaluating at $y_{i}=A^{\top}x_{\sigma(i)}+\mu$ gives
$\nabla_{y}\log f^{(\mu,A)}(y_{i})=A^{-1}\nabla\log f(x_{\sigma(i)}).$
@@
$S(Y)=\begin{pmatrix}\nabla_{y}\log f^{(\mu,A)}(y_{1})\\[-2.0pt] \vdots\\[-2.0pt] \nabla_{y}\log f^{(\mu,A)}(y_{n})\end{pmatrix}=\begin{pmatrix}A^{-1}\nabla\log f(x_{\sigma(1)})\\[-2.0pt] \vdots\\[-2.0pt] A^{-1}\nabla\log f(x_{\sigma(n)})\end{pmatrix}=P\,A^{-1}\,S(X),$
@@
Proposition A.2 (Attention can compute KDE).
@@ Appendix B Attention visualization
To interpret the outputs of the Transformer model we visualize the attention of the eight individual heads in layer 0 in two different ways. First, in Figure 12 we show the full attention matrices as heatmaps. We choose the particle ordering so that nearby particles are close in the ordering. Second, in Figure 13 we choose a random query point, marked with a red x, and color other points according to the strength of the query point’s attention. We clearly observe an emergent behavior – heads specialize in different tasks. Head 1 specialized to look at far-away points, whereas Heads 0, 2, and 5 specialized to look at close- and mid-range interactions. Finally, Heads 3, 4, 6, and 7 specialized to look in specific directions. This emergent behavior further links the multi-head Transformer to kernel-based methods, but with multi-scale learned kernels.
@@ Appendix C Comparison with the Score Matching Loss
which corresponds to the Fisher divergence $\mathbb{E}_{f}[\|s_{\theta}(x)-\nabla_{x}\log f(x)\|^{2}]$ up to a constant.
@@
Appendix D Runtime comparison
Figure 15: Runtime comparison between KDE and the Transformer model in 2 and 10 dimensions. Both are $O(n^{2})$ asymptotically, but empirically the Transformer scales better and has improved memory efficiency. KDE encounters an OOM error at $n=2^{15}$.
While KDE is known to fail in moderate and high dimensions, it is thought to be very fast. We compare the wall-clock runtime of KDE score estimation as in equation (10), and the Transformer model we propose. The Transformer model contains $800,000$ learned parameters. We use both KDE and the Transformer model on a single L40S GPU with 48Gb of memory. Both methods are $O(n^{2})$ asymptotically due to pairwise computations (kernel evaluations for KDE, attention for Transformers). However, in practice the observed scaling differs: Figure 15 shows that while KDE is faster for small sample size $n\leq 2048$, it becomes slower after this point, especially in higher dimension. Additionally, the naive KDE implementation runs out of GPU memory after $n=32,768$, while the Transformer benefits from the highly optimized attention kernels in modern deep learning frameworks, which are difficult to match with custom KDE implementations in practice.